OpenStudy (anonymous):

Suppose you divide a polynomial by a binomial. How do you know if the binomial is a factor of the polynomial? Create a sample problem that has a binomial which IS a factor of the polynomial being divided, and another problem that has a binomial which is NOT a factor of the polynomial being divided. PLEASE HELP I DO NOT UNDERSTAND THIS

OpenStudy (anonymous):

can you please give me a sample problem

OpenStudy (anonymous):

@Luis_Rivera

OpenStudy (campbell_st):

just use the factor theorem if (x -a) if a factor ot P(x) a polynomial then P(a) = 0 so by substituting a into the polynomial you will have a zero remainder

OpenStudy (campbell_st):

so if you polynomial is \[P(x) = x^2 - 5x + 6 \] to show (x -3) is factor subsitute x = 3 into P(x) \[P(3) = 3^2 - 5\times 3 + 6 = 9 = 15 + 6 = 0\] since P(3) = 0 then (x - 3) is a factor

OpenStudy (anonymous):

awesome this is great thanks

OpenStudy (campbell_st):

so for the 2nd part of the problem show (x -1) isn't a factor by finding P(1) since its not zero... then its not a factor